At the weekend I walked 2km. My dad walked 500m further than me. How far did he walk ?
Answer by lenny460(1073)

At the weekend I walked 2km. My dad walked 500m further than me. How far did he walk ?
1 km = 1000 meters
2 km = 1000 * 2 = 2000 meters
Your dad walked a further 500 meters.
That is:
2000 + 500 = 2500 meters.
Or: 2.5 km


Lennox Obuong
Algebra Tutor